# Onsager-Machlup Functional and Large Deviation Principle for Stochastic Hamiltonian Systems

### v1

@@ 5. Large Deviation Principle and Rate Function for Stochastic Hamiltonian Systems
(5.6) $\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{2}\right\}\big{|}{\left\|W^{\sigma}\right\|\leq 1}\right)=1.$
For the third term $B_{3}$,
$B_{3}$ $=\frac{1}{2}\int_{0}^{1}\left\|\sigma_{q}^{-1}(s)\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right\|^{2}\,{\rm d}s-\frac{1}{2}\int_{0}^{1}\left\|\sigma_{q}^{-1}(s)\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})\right\|^{2}\,{\rm d}s$
$\leq\frac{1}{2}\int_{0}^{1}{\sigma_{q}^{-2}(s)\left\|\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})-\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})\right\|^{2}}$
$\quad+{2\sigma_{q}^{-2}(s)\left\|\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})-\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})\right\|\left\|\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})\right\|}\,{\rm d}s$
$\leq\frac{1}{2}\int_{0}^{1}{\sigma_{q}^{-2}(s)\left\|\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})-\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})\right\|^{2}}\,{\rm d}s$
$\quad+\int_{0}^{1}{\sigma_{q}^{-2}(s)\left\|\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})-\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})\right\|\left\|\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})\right\|}\,{\rm d}s.$
Using that $\frac{\partial H}{\partial p}$ is Lipschitz continuous, we have
(5.7) $\quad\left\|\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})-\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right\|$
$=\left\|\frac{\partial H}{\partial\left(\varphi_{p}+\epsilon W^{\sigma}_{p}\right)}((\varphi_{q}+\epsilon W^{\sigma}_{q}),(\varphi_{p}+\epsilon W^{\sigma}_{p}))-\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right\|$
$\leq L\epsilon\left\|W^{\sigma}\right\|.$
[…]
(5.8) $\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{3}\right\}\big{|}{\left\|W^{\sigma}\right\|\leq 1}\right)=1$
@@
For the fourth term $B_{4}$, employing the same proof technique as for the third term $B_{3}$, we have
(5.9) $\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{4}\right\}\big{|}{\left\|W^{\sigma}\right\|\leq 1}\right)=1$
for all $c\in\mathbb{R}$.
[…]
$B_{5}$ $=\int_{0}^{1}\left\langle\sigma_{q}^{-2}(s)\left(\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})-\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right),\dot{\varphi}_{q}(s)\right\rangle\,{\rm d}s$
$\leq C\left\|\frac{\partial H}{\partial y_{p}}(y_{q},y_{p})-\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right\|$
$\leq CL\epsilon\left\|W^{\sigma}\right\|.$
Thus,
(5.10) $\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{5}\right\}\big{|}{\left\|W^{\sigma}\right\|\leq 1}\right)=1$
for all $c\in\mathbb{R}$.
For the sixth term $B_{6}$, employing the same proof technique as for the fifth term $B_{5}$, we have
(5.11) $\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{6}\right\}\big{|}{\left\|W^{\sigma}\right\|\leq 1}\right)=1$
@@
For each compact set $K_{n}$ and any $\epsilon>0$, there always exist a finite number of continuous curves $\{\varphi_{j}\}_{j=1}^{N}\subset K_{n}$ and their corresponding neighborhoods $K(\varphi_{j},\epsilon)$ such that:
@@
Applying the union bound principle to this finite covering, we obtain:
@@
Taking the logarithm of the above inequality and multiplying by $\epsilon^{2}$, we get:
@@
Taking the upper limit $\limsup\limits_{\epsilon\to 0}$ of the above equation, we obtain:
@@
Since $\epsilon^{2}\ln N$ tends to 0, the dominant term is determined by the neighborhood with the maximum probability. Combining the above proofs, we have:
$\limsup_{\epsilon\to 0}\epsilon^{2}\ln\mathbb{P}((X(t)\in K_{n})$ $\leq\lim_{\epsilon\to 0}\max_{1\leq j\leq N}\left(-\frac{1}{2}\int_{0}^{1}U(\varphi_{j})\,{\rm d}s\right.$
$\quad\left.+\mathbb{E}\left(\sum_{i=1}^{6}B_{i}\mid\left\|W^{\sigma}\right\|\leq 1\right)+\epsilon^{2}\ln\mathbb{P}(\left\|W^{\sigma}\right\|\leq 1)\right)$
@@
Since $\{\varphi_{j}\}\subset K_{n}$, we get:
@@
For any $\delta>0$, there exists a compact set $K_{n}\subset F$ such that:
@@
Since $\mathbb{P}((X(t)\in F)\leq\mathbb{P}((X(t)\in K_{n})+\mathbb{P}((X(t)\in F\setminus K_{n})$, and the probability of the part outside the compact set decays exponentially to 0 as $n$ increases, the main contribution comes from $K_{n}$. Therefore, as $n$ tends to infinity and $\epsilon$ tends to 0, we obtain:
(5.12) $\limsup_{\epsilon\to 0}\epsilon^{2}\ln\mathbb{P}((X(t)\in F)\leq-\inf_{\varphi\in F}\frac{1}{2}\int_{0}^{1}U(\varphi_{j})\,{\rm d}s.$
@@
(5.13) $\liminf_{{\epsilon\to 0}}\epsilon^{2}\ln\mathbb{P}(X(t)\in G)\geq-\inf_{\varphi\in G}\frac{1}{2}\int_{0}^{1}U(\varphi)\,{\rm d}s.$
By combining the upper bound inequality (5.12) and the lower bound inequality (5.13), we can derive the rate function for the stochastic Hamiltonian system:
$J(\varphi)$ $=\frac{1}{2}\left(\int_{0}^{1}\left\|\sigma_{q}^{-1}(t)\left(\dot{\varphi}_{q}-\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right)\right\|^{2}\,{\rm d}t\right.$
$\qquad\left.+\int_{0}^{1}\left\|\sigma_{p}^{-1}(t)\left(\dot{\varphi}_{p}+\frac{\partial H}{\partial\varphi_{q}}(\varphi_{q},\varphi_{p})\right)\right\|^{2}\,{\rm d}t\right).$
6. Preservation of Invariant Tori in Nearly Integrable Stochastic Hamiltonian Systems
@@
  • The Lebesgue (outer) measure on $\mathbb{R}^{d}$ is denoted by $meas$.
@@
  • For $l>0$ and an open subset $A$ of $\mathbb{R}^{d}$ or $\mathbb{R}^{d}\times\mathbb{T}^{d}$, the set $C^{l}(A)$ consists of continuously differentiable functions $f$ on $A$ up to the order $[l]$ such that $f^{[l]}$ is Hölder-continuous with exponent $\{l\}$ and with finite $C^{l}$-norm defined by:
    $\|f\|_{C^{l}(A)}$ $:=\max\left\{\|f\|_{C^{[l]}(A)},\|f^{[l]}\|_{C^{\{l\}}(A)}\right\},$
    $\|f\|_{C^{[l]}(A)}$ $:=\max_{\begin{subarray}{c}k\in\mathbb{N}^{d}\\ 0\leq|k|_{1}\leq[l]\end{subarray}}\sup_{A}|\partial_{I}^{k}f|,$
    $\|f^{[l]}\|_{C^{\{l\}}(A)}$ $:=\max_{\begin{subarray}{c}k\in\mathbb{N}^{d}\\ |k|_{1}=[l]\end{subarray}}\sup_{\begin{subarray}{c}I_{1},I_{2}\in A\\ 0<|I_{1}-I_{2}|<1\end{subarray}}\frac{|\partial_{I}^{k}f(I_{1})-\partial_{I}^{k}f(I_{2})|}{|I_{1}-I_{2}|^{\{l\}}}.$
@@
  • For $r,s>0$, $y_{0}\in\mathbb{C}^{d}$, and $\varnothing\neq\mathscr{D}\subseteq\mathbb{C}^{d}$, we define:
@@
[…]
    where $e:=\exp(1)$ denotes the Neper number and $i$ the imaginary unit. We also set:
[…]
Without loss of generality, the Hamiltonian equations for an integrable system are expressed as:
(6.1) $\begin{cases}d\theta(t)=\frac{\partial H}{\partial I}(I(t))\,dt=\omega(I(t))\,dt,\\ dI(t)=0,\end{cases}$
@@
When introducing a small deterministic perturbation $\epsilon_{1}P(I,\theta)$, the almost integrable Hamiltonian becomes:
@@
(6.2) $\begin{cases}\mathrm{d}\theta(t)=\frac{\partial H_{\epsilon}}{\partial I}(I(t),\theta(t))\,\mathrm{d}t,\\ \mathrm{d}I(t)=-\frac{\partial H_{\epsilon}}{\partial\theta}(I(t),\theta(t))\,\mathrm{d}t.\end{cases}$
@@
(6.3) $\begin{cases}\mathrm{d}\theta(t)=\frac{\partial H_{\epsilon}}{\partial I}(I(t),\theta(t))\,\mathrm{d}t+\epsilon_{2}\sigma_{\theta}(t)\,\mathrm{d}W_{\theta}(t),\\ \mathrm{d}I(t)=-\frac{\partial H_{\epsilon}}{\partial\theta}(I(t),\theta(t))\,\mathrm{d}t+\epsilon_{2}\sigma_{I}(t)\,\mathrm{d}W_{I}(t),\end{cases}$
where $\epsilon_{2}$ represents the strength of the stochastic perturbation, $\sigma_{\theta}(t)$ and $\sigma_{I}(t)$ are the diffusion coefficients, and $W_{\theta}(t)$ and $W_{I}(t)$ are independent Wiener processes.
We then investigate the invariant tori associated with the Hamiltonian
$H_{\epsilon}(I,\theta)=H(I)+\epsilon_{1}P(I,\theta),$
corresponding to Equation (6.2). The following assumptions are imposed:
@@
    Consider the Hamiltonian $H_{\epsilon}(I,\theta)$ on the phase space $\mathscr{D}\times\mathbb{T}^{d}$. Here, $H$ and $P$ are given functions in $C^{l}(\mathscr{D}\times\mathbb{T}^{d})$ with finite $l$ norms $\|H\|_{C^{l}(\mathscr{D})}$ and $\|P\|_{C^{l}(\mathscr{D})}$.
@@
$OM(\varphi_{\theta},\varphi_{I})$ $=\frac{1}{\epsilon_{2}^{2}}\left(\int_{0}^{1}\left\|\sigma_{\theta}^{-1}(t)\left(\dot{\varphi_{\theta}}-\omega(\varphi_{I})-\epsilon_{1}\frac{\partial P}{\partial\varphi_{I}}(\varphi_{\theta},\varphi_{I})\right)\right\|^{2}\,{\rm d}t\right.$
$\qquad\left.+\int_{0}^{1}\left\|\sigma_{I}^{-1}(t)\left(\dot{\varphi_{I}}+\epsilon_{1}\frac{\partial P}{\partial\varphi_{\theta}}(\varphi_{\theta},\varphi_{I})\right)\right\|^{2}\,{\rm d}t\right).$
@@
[…]
$\epsilon_{2}^{2}\ln\mathbb{P}(X_{\epsilon}(t)\in A)\approx-\inf_{\varphi\in A}I(\varphi),$
where $A\in\mathbb{R}^{d}$ denotes an arbitrary measurable set and the rate function $I(\varphi)$ is given by:
$I(\varphi)$ $=\frac{1}{2}\left(\int_{0}^{1}\left\|\sigma_{\theta}^{-1}(t)\left(\dot{\varphi_{\theta}}-\omega(\varphi_{I})-\epsilon_{1}\frac{\partial P}{\partial\varphi_{I}}(\varphi_{\theta},\varphi_{I})\right)\right\|^{2}\,{\rm d}t\right.$
$\qquad\left.+\int_{0}^{1}\left\|\sigma_{I}^{-1}(t)\left(\dot{\varphi_{I}}+\epsilon_{1}\frac{\partial P}{\partial\varphi_{\theta}}(\varphi_{\theta},\varphi_{I})\right)\right\|^{2}\,{\rm d}t\right).$
@@
Defining $\tilde{H}:=\widehat{H}+\chi\cdot(H-\widehat{H})$, we ensure $\tilde{H}\in C^{l}(\mathbb{R}^{d}\times\mathbb{T}^{d})$ and $\tilde{H}\equiv H$ on $D_{\alpha_{*}/2}(\mathscr{D}^{\prime})$. Furthermore,
@@
where $R(\mathcal{D}):=\sup\{R>0:B_{R}(I)\subseteq\mathcal{D}\text{ for some }I\in\mathcal{D}\}$, $\mathrm{minfoc}(\partial\mathcal{D})$ is the minimal focal distance of $\partial\mathcal{D}$ and $a:=(l-2v)^{-1}\max\left\{(6+2lv^{-1})(l+v)-2l(l-v),\right.$ $\left.2l(l+3v)v^{-1}\right\}$.
Let $\mathcal{H}_{j}$ (resp. $\mathcal{P}_{j}$) denote the real-analytic approximation $H_{\xi_{j}}$ (resp. $P_{\xi_{j}}$) of $\tilde{H}$ (resp. $\tilde{P}$) defined on $\mathcal{O}_{j}$, as given by Lemma 2.15. Define the initial set $\mathscr{D}_{0}$ as:
@@
$\left\lVert\partial_{I}^{2}H_{j}\right\rVert_{r_{j},\mathscr{D}_{j}}<2C_{H},\quad\left\lVert T_{j}\right\rVert_{\mathscr{D}_{j}}<2C_{T},\quad T_{j}:=(\partial_{I}^{2}H_{j})^{-1},\quad\left\lVert P_{j}\right\rVert_{r_{j},s_{j},\mathscr{D}_{j}}\leq C_{1}C_{H}\xi_{j}^{l-1},$
@@
(6.4) $\phi_{j}(D_{r_{j},s_{j}}(\mathscr{D}_{j}))\subset D_{\sigma_{j-1},\sigma_{j-1}}(\mathscr{D}_{j-1})\bigcap D_{r_{j-1}/2,s_{j-1}/2}(\mathscr{D}_{j-1}).$
@@
holds. Together with the definitions of the sequences of the various parameters, this ensures that condition (2.2) in Lemma 2.14 is satisfied for all $j\geq 1$.
@@
$\mathcal{H}^{\epsilon}_{j}:=\mathcal{H}_{j}+\mathcal{P}_{j}=\mathcal{H}^{\epsilon}_{j-1}+(\mathcal{H}_{j}-\mathcal{H}_{j-1})+(\mathcal{P}_{j}-\mathcal{P}_{j-1})$
By the inductive assumption and (6.4), we have
@@
(6.5) $\left\lVert\partial_{y}^{2}\mathcal{H}^{j}\right\rVert_{r_{j},\mathscr{D}_{j}}<2C_{K},\qquad\left\lVert(\partial_{y}^{2}\mathcal{H}^{j})^{-1}\right\rVert_{\mathscr{D}_{j}}<2C_{T},$

### v2

@@ 5 Large Deviation Principle and Rate Function for Stochastic Hamiltonian Systems
$\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{2}\right\}\big{|}{\left\|W^{\sigma}\right\|\leq 1}\right)=1.$ (34)
For the third term $B_{3}$, let $v:=J\nabla H(y_{s})-\dot{\varphi}_{s}$, then
$B_{3}$ $=\frac{1}{2}\int_{0}^{1}\left(\left|\sigma^{-1}(s,\varphi_{s})v\right|^{2}-\left|\sigma^{-1}(s,y_{s})v\right|^{2}\right)\,{\rm d}s$
$\leq\frac{1}{2}\int_{0}^{1}{\left(\left\|\left({\sigma^{-1}\left(s,y_{s}\right)}\right)^{2}-\left({\sigma^{-1}\left(s,\varphi_{s}\right)}\right)^{2}\right\|_{HS}\right)\left|v\right|^{2}}\,{\rm d}s$
$\leq\frac{1}{2}\int_{0}^{1}{\left\|{\sigma^{-1}\left(s,y_{s}\right)}-{\sigma^{-1}\left(s,\varphi_{s}\right)}\right\|_{2}\left\|{\sigma^{-1}\left(s,y_{s}\right)}+{\sigma^{-1}\left(s,\varphi_{s}\right)}\right\|_{2}\left|v\right|^{2}}\,{\rm d}s$
$\leq\frac{1}{2}\int_{0}^{1}\left\|{\sigma^{-1}\left(s,y_{s}\right)}\right\|_{2}\left\|{\sigma\left(s,\varphi_{s}\right)-\sigma\left(s,y_{s}\right)}\right\|_{2}\left\|{\sigma^{-1}\left(s,\varphi_{s}\right)}\right\|_{2}$
$\qquad\times\left\|{\sigma^{-1}\left(s,y_{s}\right)}+{\sigma^{-1}\left(s,\varphi_{s}\right)}\right\|_{2}\left|v\right|^{2}\,{\rm d}s.$
Since both $\sigma^{-1}$ and $v$ are bounded, and $\sigma$ is Lipschitz continuous, while $\varphi_{t}-X(0)\in\mathbb{H}^{1}$, we conclude that there exists a constant $C>0$ such that
$B_{3}\leq C\left\|\sigma(s,\varphi_{s})-\sigma(s,y_{s})\right\|_{2}\leq\epsilon CL\left\|W^{\sigma}\right\|.$
The above inequality implies that
$\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{3}\right\}\big{|}\|W^{\sigma}\|\leq 1\right)=1$ (35)
@@
For the fourth term $B_{4}$, let $v:=J\nabla H(y_{s})-\dot{\varphi}_{s}$, $u:=J\nabla H(\varphi_{s})-\dot{\varphi}_{s}$, then.
$B_{4}$ $=\frac{1}{2}\int_{0}^{1}{\left|\sigma^{-1}(s,\varphi_{s})u\right|^{2}-\left|\sigma^{-1}(s,\varphi_{s})v\right|^{2}}\,{\rm d}s$
$=\frac{1}{2}\int_{0}^{1}{\left\langle\sigma^{-1}(s,\varphi_{s})^{2}\left(u-v\right),\left(u+v\right)\right\rangle}\,{\rm d}s$
$\leq\frac{1}{2}\int_{0}^{1}{\left\|\sigma^{-1}(s,\varphi_{s})^{2}\right\|_{2}\left|\left(u-v\right)\right|\left|\left(u+v\right)\right|}\,{\rm d}s.$
Since both $\sigma^{-1}$, $v$, and $u$ are bounded, and $J\nabla H(y_{s})$ is Lipschitz continuous, and noting that $\varphi_{t}-X(0)\in\mathbb{H}^{1}$, we can conclude that there exists a constant $C>0$ such that
$B_{4}\leq C\left|u-v\right|\leq C\left|J\nabla H(\varphi_{s})-J\nabla H(y_{s})\right|\leq\epsilon CL\left\|W^{\sigma}\right\|.$
The above inequality implies that
$\limsup\limits_{\epsilon\to 0}\mathbb{E}\left({\rm exp}\left\{cB_{4}\right\}\big{|}\|W^{\sigma}\|\leq 1\right)=1$ (36)
@@
For each compact set $K_{n}$ and any $\epsilon>0$, there always exist a finite number of continuous curves $\{\varphi_{j}\}_{j=1}^{N}\subset K_{n}$ and their corresponding neighborhoods $K(\varphi_{j},\epsilon)$ such that
@@
Applying the union bound principle to this finite covering, we obtain
@@
Taking the logarithm of the above inequality and multiplying by $\epsilon^{2}$, we get
@@
Taking the upper limit $\limsup\limits_{\epsilon\to 0}$ of the above equation, we obtain
@@
Since $\epsilon^{2}\ln N$ tends to 0, the dominant term is determined by the neighborhood with the maximum probability. Combining the above proofs, we have
$\quad\limsup_{\epsilon\to 0}\epsilon^{2}\ln\mathbb{P}((X(t)\in K_{n})$
$\leq\lim_{\epsilon\to 0}\max_{1\leq j\leq N}\left(-\frac{1}{2}\int_{0}^{1}U(\varphi_{j})\,{\rm d}s+\mathbb{E}\left(\sum_{i=1}^{6}B_{i}\mid\left\|W^{\sigma}\right\|\leq 1\right)+\epsilon^{2}\ln\mathbb{P}(\left\|W^{\sigma}\right\|\leq 1)\right)$
@@
Since $\{\varphi_{j}\}\subset K_{n}$, we get
@@
For any $\delta>0$, there exists a compact set $K_{n}\subset F$ such that
@@
Since $\mathbb{P}((X(t)\in F)\leq\mathbb{P}((X(t)\in K_{n})+\mathbb{P}((X(t)\in F\setminus K_{n})$, and the probability of the part outside the compact set decays exponentially to 0 as $n$ increases, the main contribution comes from $K_{n}$. Therefore, as $n$ tends to infinity and $\epsilon$ tends to 0, we obtain
$\limsup_{\epsilon\to 0}\epsilon^{2}\ln\mathbb{P}((X(t)\in F)\leq-\inf_{\varphi\in F}\frac{1}{2}\int_{0}^{1}U(\varphi_{j})\,{\rm d}s.$ (37)
@@
$\liminf_{{\epsilon\to 0}}\epsilon^{2}\ln\mathbb{P}(X(t)\in G)\geq-\inf_{\varphi\in G}\frac{1}{2}\int_{0}^{1}U(\varphi)\,{\rm d}s.$ (38)
By combining the upper bound inequality (37) and the lower bound inequality (38), we can derive the rate function for the stochastic Hamiltonian system
$S(\varphi)=\frac{1}{2}\left(\int_{0}^{1}\left\|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi_{s}}\right)\right\|^{2}\,{\rm d}s\right).$
6 Preservation of Invariant Tori in Nearly Integrable Stochastic Hamiltonian Systems
@@
  • The Lebesgue (outer) measure on $\mathbb{R}^{d}$ is denoted by measmeas.
@@
  • For $l>0$ and an open subset $A$ of $\mathbb{R}^{d}$ or $\mathbb{R}^{d}\times\mathbb{T}^{d}$, the set $C^{l}(A)$ consists of continuously differentiable functions $f$ on $A$ up to the order $[l]$ such that $f^{[l]}$ is Hölder-continuous with exponent $\{l\}$ and with finite $C^{l}$-norm defined by
    $\|f\|_{C^{l}(A)}$ $=\max\left\{\|f\|_{C^{[l]}(A)},\|f^{[l]}\|_{C^{\{l\}}(A)}\right\},$
    $\|f\|_{C^{[l]}(A)}$ $=\max_{\begin{subarray}{c}k\in\mathbb{N}^{d}\\ 0\leq|k|_{1}\leq[l]\end{subarray}}\sup_{A}|\partial_{I}^{k}f|,$
    $\|f^{[l]}\|_{C^{\{l\}}(A)}$ $=\max_{\begin{subarray}{c}k\in\mathbb{N}^{d}\\ |k|_{1}=[l]\end{subarray}}\sup_{\begin{subarray}{c}I_{1},I_{2}\in A\\ 0<|I_{1}-I_{2}|<1\end{subarray}}\frac{|\partial_{I}^{k}f(I_{1})-\partial_{I}^{k}f(I_{2})|}{|I_{1}-I_{2}|^{\{l\}}}.$
@@
  • For $r,s>0$, $y_{0}\in\mathbb{C}^{d}$, and $\varnothing\neq\mathscr{D}\subseteq\mathbb{C}^{d}$, we define
@@
    $f=\sum_{k\in\mathbb{Z}^{d}}f_{k}e^{ik\cdot\theta},\quad f_{k}:=\frac{1}{(2\pi)^{d}}\int_{\mathbb{T}^{d}}f(\theta)e^{-ik\cdot\theta}d\theta,$
    where $e:=\exp(1)$ denotes the Neper number and $i$ the imaginary unit. We also set
    $\langle f\rangle:=f_{0}=\frac{1}{(2\pi)^{d}}\int_{\mathbb{T}^{d}}f(\theta)d\theta.$
Without loss of generality, the Hamiltonian equations for an integrable system are expressed as
$\begin{cases}d\theta(t)=\frac{\partial H}{\partial I}(I(t))\,dt=\omega(I(t))\,dt,\\ dI(t)=0,\end{cases}$ (39)
@@
When introducing a small deterministic perturbation $\epsilon_{1}P(I,\theta)$, the almost integrable Hamiltonian becomes
@@
$\mathrm{d}X(t)=J\nabla H_{\epsilon}(X(t))\,\mathrm{d}t.$ (40)
@@
$\mathrm{d}X(t)=J\nabla H_{\epsilon}(X(t))\,\mathrm{d}t+\epsilon_{2}\sigma(t,X(t))\,\mathrm{d}W(t),$ (41)
where $\epsilon_{2}$ represents the strength of the stochastic perturbation, $\sigma(t,X(t))$ denotes the diffusion coefficient matrix, and $W(t)$ is a $2n$-dimensional Wiener process with independent components.
We then investigate the invariant tori associated with the Hamiltonian $H_{\epsilon}(I,\theta)$ corresponding to Equation (40). The following assumptions are imposed:
@@
    Consider the Hamiltonian $H_{\epsilon}(I,\theta)$ on the phase space $\mathscr{D}\times\mathbb{T}^{d}$. Here, $H$ and $P$ are given functions in $C^{l}(\mathscr{D}\times\mathbb{T}^{d})$ with finite $l$ - norms $\|H\|_{C^{l}(\mathscr{D})}$ and $\|P\|_{C^{l}(\mathscr{D})}$.
@@
$OM(\varphi,\dot{\varphi})$ $=\int_{0}^{1}{\left|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H_{\epsilon}(\varphi_{s})-\dot{\varphi_{s}}\right)\right|^{2}}\,{\rm d}s$
$\quad-\int_{0}^{1}\mathrm{Tr}\left(\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\left(J\nabla H_{\epsilon}(\varphi_{s})-\dot{\varphi}_{s}\right)\right)\,{\rm d}s$
Furthermore, by applying the variational principle to minimize the Onsager-Machlup functional, the most probable transition path can be obtained, corresponding to the solution of the nearly integrable Hamiltonian system
$\dot{\varphi}(t)=J\nabla H_{\epsilon}(\varphi_{t})-\frac{1}{2}\epsilon_{2}^{2}\sigma(t,\varphi_{t})\sum_{i}\left(D_{y}\sigma(t,\varphi_{t})\right)_{\cdot ii}.$
@@
$\epsilon_{2}^{2}\ln\mathbb{P}(X_{\epsilon}(t)\in A)\approx-\inf_{\varphi\in A}S(\varphi),$
where $A\in\mathbb{R}^{d}$ denotes an arbitrary measurable set and the rate function $S(\varphi)$ is given by
$S(\varphi)=\frac{1}{2}\left(\int_{0}^{1}\left\|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H_{\epsilon}(\varphi_{s})-\dot{\varphi_{s}}\right)\right\|^{2}\,{\rm d}s\right).$
@@
[…]
@@
[…]
Let $\mathcal{H}_{j}$ (resp. $\mathcal{P}_{j}$) denote the real-analytic approximation $H_{\xi_{j}}$ (resp. $P_{\xi_{j}}$) of $\tilde{H}$ (resp. $\tilde{P}$) defined on $\mathcal{O}_{j}$, as given by Lemma 2.17. Define the initial set $\mathscr{D}_{0}$ as
@@
$\left\lVert\partial_{I}^{2}H_{j}\right\rVert_{r_{j},\mathscr{D}_{j}}<2C_{H},\quad\left\lVert T_{j}\right\rVert_{\mathscr{D}_{j}}<2C_{T},$
$T_{j}:=(\partial_{I}^{2}H_{j})^{-1},\quad\left\lVert P_{j}\right\rVert_{r_{j},s_{j},\mathscr{D}_{j}}\leq C_{1}C_{H}\xi_{j}^{l-1},$
@@
$\phi_{j}(D_{r_{j},s_{j}}(\mathscr{D}_{j}))\subset D_{\sigma_{j-1},\sigma_{j-1}}(\mathscr{D}_{j-1})\bigcap D_{r_{j-1}/2,s_{j-1}/2}(\mathscr{D}_{j-1}).$ (42)
@@
holds. Together with the definitions of the sequences of the various parameters, this ensures that condition (4) in Lemma 2.16 is satisfied for all $j\geq 1$.
@@
$\mathcal{H}^{\epsilon}_{j}:=\mathcal{H}_{j}+\mathcal{P}_{j}=\mathcal{H}^{\epsilon}_{j-1}+(\mathcal{H}_{j}-\mathcal{H}_{j-1})+(\mathcal{P}_{j}-\mathcal{P}_{j-1}).$
By the inductive assumption and (42), we have
@@
$\left\lVert\partial_{y}^{2}\mathcal{H}^{j}\right\rVert_{r_{j},\mathscr{D}_{j}}<2C_{K},\qquad\left\lVert(\partial_{y}^{2}\mathcal{H}^{j})^{-1}\right\rVert_{\mathscr{D}_{j}}<2C_{T},$ (43)